Suppose Sashac, a champion bicyclist, wants to see how far he can travel in an hour. He stats timing himself when he reaches a speed of 45 miles/hour. He maintains this speed for 10 minutes. Sascha starts to feel tired and slows down to 30 m/h for the next 5 minutes. He then reduces his speed to 25 m/h for the next 30 minutes. Finally, Sascha feels exhausted as he finishes the last 15 minutes at 15 m / h. I get the average speed he travels is 28.75 but get that he travels only 26.25 and if he travels for an hour it doesn't make sense what did I do wrong. Thanks: Suppose Sashac, a champion bicyclist, wants to see how far he can travel in an hour. He stats timing himself when he reaches a speed of 45 miles/hour. He maintains this speed for 10 minutes. Sascha starts to feel tired and slows down to 30 m/h for the next 5 minutes. He then reduces his speed to 25 m/h for the next 30 minutes. Finally, Sascha feels exhausted as he finishes the last 15 minutes at 15 m / h. I get the average speed he travels is 28.75 m/h but get that he travels only 26.25 when using d=(r)(t)and if he travels for an hour it doesn't make sense what did I do wrong. Thanks
Answer by scott8148(6628)

not sure how you found average speed but finding the total distance is the best method

45(2/12)+30(1/12)+25(6/12)+15(3/12)=(90+30+150+45)/12=315/12=105/4=26.25

the question only asks for distance

the average speed would be a weighted average because the different speeds last for different times
___ this could be where you went astray